OpenStudy (aaaa):

Write the equation y - 2 = 2(x - 3) in standard form. A. y - 2x = 4 B. 2x - y = -6 C. 2x + y = 6 D. 2x - y = 4

OpenStudy (ash2326):

y-2=2x-6 2x-y=-4 or y-2x=4 option A